IPP Progress Report for the week 2010.08.23 - 2010.08.27

(Up to IPP Progress Reports)

Eugene Magnier

Serge Chastel

* PSPS:

  • Got some input files from Roy
  • Started to implement the tests. I'm also detailing the changes for the IPP-PSPS ICD.

* IPP:

  • IPP compilation test for Hudson is now ok (I still have small problems with dbconfig/ippdb though). All IPP modules are registered. Dependencies between them need to be checked.

* Made a talk about Hudson for Dev's Lunch.

Heather Flewelling

Roy Henderson

* PSPS

  • tested Conrad's DXLayer changes on sandbox
  • resumed loading to Maui after big merge over the weekend. Plan to work though all 3PI available in DVO Db

* IPP

  • finally wrote XML schema for the BatchManifest files for ippToPsps. Serge can use this as part of his testing.
  • czartool
    • wrote functions to trim czar Db so that it has data per hour rather than per minute. Most recent days will keep higher resolution.
    • new code and a Db table to log burntool progress
    • new code and Db table to store current nightly science status
    • Changed czartool page to included new burntool data and nightly science status
  • czar on Friday

Bill Sweeney

  • Installed new version of IPP at MPIA. Ran into some build problems with optimized code on old compiler. (This is a known issue but I forgot).
  • Spent several hours debugging some issues with dvomerge for MPIA. We now have a working method for merging a subset of ps1 and sdss dvo catalogs
  • debugged problem with the "lossy compression" system
  • Accidentally deleted the reference stack bundles from the distribution data store again. Need to find a way to prevent this from happening.
  • Spent two days as processing czar. I get little useful work done on these days.
  • Ran into space crunch again. Spent far too much time trying to recover space and dealing with flaky nodes. Aren't computers and disks cheaper than me?

Chris Waters

  • Linearity: continued work to implement software fix in ppImage.
  • Disk Space: tried to figure out where all our disk space has gone. Tracked at least some of it down to over-duplicated data. Have implemented a neb-countcheck program that can be used to scan the database for data that either needs to be culled or replicated, but have not added the pantasks back end that is needed to do this scanning.
  • Detectability: fixed bug in update processing that should allow requests to process correctly. Received email about a post-fix request that was processing correctly.
  • Shutter mask: looked at a few hundred exposures to confirm that the mask seems to block out star glints.
  • Liason: queued up requested data for JHU. For this data, data_group LIKE '%czw%';
  • Bug fixes for new tag (psLib, psModules, ippmonitor).
  • Flying to Belfast.